What is the value of k to 3 significant figures (include UNITS) for the reaction with the rate law R = k[A]1[B]2 if the rate is 0.0450M/s, [A]=0.100M and [B]=0.200M?


Show Your Work


1
Expert's answer
2021-01-11T04:45:08-0500

R = k[A][B]2

0.0450 = k(0.100 * 0.2002)

k = 0.0450/(0.1 *0.04)

k = 0.0450/0.004

k = 11.25